# ReDiFine: Reusable Diffusion Finetuning for Mitigating Degradation in the Chain of Diffusion

### v1

@@ Abstract
Diffusion models have achieved tremendous improvements in generative modeling for images, enabling high-quality generation that is indistinguishable by humans from real images. The qualities of images have reached a threshold at which we can reuse synthetic images for training machine learning models again. This attracts the area as it can relieve the high cost of data collection and fundamentally solve many problems in data-limited areas. In this paper, we focus on a practical scenario in which pretrained text-to-image diffusion models are iteratively finetuned using a set of synthetic images, which we call the Chain of Diffusion. Finetuned models generate images that are used for the next iteration of finetuning. We first demonstrate how these iterative processes result in severe degradation in image qualities. Thorough investigations reveal the most impactful factor for the degradation, and we propose finetuning and generation strategies that can effectively resolve the degradation. Our method, Reusable Diffusion Finetuning (ReDiFine), combines condition drop finetuning and CFG scheduling to maintain the qualities of generated images throughout iterations. ReDiFine works effectively for multiple datasets and models without further hyperparameter search, making synthetic images reusable to finetune future generative models.
@@ 1 Introduction
Can state-of-the-art AI models learn from their own outputs and improve themselves? As generative AI (e.g., GPT [1], Diffusion [2, 3]) can now churn out uncountable synthetic texts and images, this question piqued curiosity from many researchers in the past couple of years [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. While some show positive results of self-improving [4, 15], most report an undesirable “model collapse” [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 16, 17, 18]—a phenomenon where a model’s performance degrades when it goes through multiple cycles of training with the self-generated data. When large language models (LLMs) are trained with their own outputs, it begins to produce low-quality text that has a lot of repetitions [19], and its linguistic diversity declines rapidly [8, 16]; image models also show quality degradation [7, 10] and loss of diversity [12, 20]. This phenomenon has even been compared to mad cow disease, a fatal neurodegenerative disease caused by feeding cattle with the carcasses of other cattle [12].
Our work examines model collapse in a common use case of generative AI: customizing pretrained text-to-image generative models through finetuning. While previous works [5, 12, 14, 19] suggest including more real images in the training set to avoid model collapse, this is challenging for several reasons. First, the number of artworks generated by a single artist is often limited. Even highly productive artists create at most a few thousands of pieces during their lifetime (e.g., Monet produced 1,983 paintings over 61 years). Secondly, during image collection, it is easy to be deceived by realistic AI-generated images and accidentally include them in the training set. Therefore, we need to understand mode collapse behavior when the proportion of real data is limited and seek solutions beyond simply increasing the number of real images. In this paper, we propose a novel concept of generating “reusable” images, which refers to their potential for finetuning without causing catastrophic model collapse.
(a) Self-Consuming Chain of Diffusion.
(b) Image degradation with the Chain of Diffusion.
Figure 1: (a) Overall pipeline of the Chain of Diffusion. Given a pretrained text-to-image diffusion model $M_{0}$ and a prompt set $P$, a finetuned model $M_{k}$ is trained using $D_{k-1}$ generated at the previous iteration $k-1$. Then, $M_{k}$ generates $D_{k}$ using the same prompt set $P$. Chain of diffusion begins with the original dataset $D_{0}$. (b) Image degradation universally occurs during Chain of Diffusion across four datasets. As the Chain of Diffusion progresses, the severity of image degradation intensifies, exhibiting universal patterns of highly saturated images containing repetitive high-frequency patterns. This consistently holds across four datasets and $10$ scenarios that we comprehensively investigate in Section 3.2.
Our goal in this paper is two-fold: (1) to conduct an in-depth empirical analysis of model collapse behavior in the self-consuming chain of diffusion finetuning (referred to as Chain of Diffusion)—a setting already widely used by artists—and to examine which factors in finetuning and sampling affect the reusability of diffusion-generated images; (2) to develop a strategy that can improve the reusability of synthetic images. Our contributions toward these goals are summarized below:
  • We consider a setting widely used for style transfer in diffusion models: finetuning Stable Diffusion [3] with low-rank adaptation (LoRA) finetuning [21] using a moderately-sized training set($\sim$1000100010001000 images). We examine the evolution of images in the Chain of Diffusion across four datasets: two digital art datasets (Pokemon [22] and Kumapi [23]) and two natural image datasets (Butterfly [24] and CelebA-1k [25]). Through extensive analysis, we show that model collapse universally occurs in all datasets regardless of training set size, real image mixing (up to 90%), and a wide range of scenarios with different configurations (see Figure 11(b) and Table 1).
  • We discover that among all factors in the Chain of Diffusion, only one substantially impacts the speed of model collapse: classifier-free guidance (CFG) scale. With low CFG, we observe low-frequency degradation where images gradually become unrecognizably blurry. With high CFG scales, high-frequency degradation occurs, where some features of the images become amplified beyond the normal range with disturbing color saturation. There exists a medium CFG range that can significantly slow down model collapse (see Figure 22(b)). To the best of our knowledge, we are the first to demonstrate such a role of CFG in the Chain of Diffusion.
  • Further, we reveal that the default CFG scale $7.5$ widely used for Stable Diffusion performs very poorly in terms of reusability metric (defined in Eqn (4)) despite its excellent image quality in the first iteration (measured by FID). This highlights the trade-off between first-iteration image quality and future reusability (see Figure 23).
  • We develop a new finetuning strategy, Reusable Diffusion Finetuning (ReDiFine), to mitigate degradation inspired by our analysis of how CFG modulates the model collapse pattern in latent diffusion models. We combine condition drop (for finetuning) and CFG scheduling (for generation) in ReDiFine. While an optimal CFG scale can sufficiently slow down model collapse, it is impractical to sweep through CFG scales over multiple iterations for each dataset to find an optimal trade-off. With our proposed ReDiFine approach, we show that even the default CFG scale (CFG=7.5) can generate reusable images across all four datasets without the need for parameter tuning (see Figure 5 and 6).
@@ 2 Related Work
The self-consuming training loop and the associated phenomenon known as “model collapse” have become significant areas of study in the past two years [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Model collapse, defined as “a degenerative process affecting generations of learned generative models, where generated data end up polluting the training set of the next generation of models” in [11], has been observed in both language and image generative models. Empirical studies on LLMs [8, 15, 16] reveal that linguistic diversity collapses, especially in high-entropy tasks [8], although this can be mitigated with data accumulation [15]. In image generation, several works [5, 7, 10, 12, 18, 20] note image degradation when diffusion models are recursively trained with self-generated data. The authors of [5] and [12] provide theoretical insights into diversity reduction, demonstrating that the covariance in a toy Gaussian model converges to zero as iterations increase in the self-consuming training loop. Among theoretical studies on this topic [13, 14, 17, 19], [17] is particularly relevant. They use dynamical system tools to show that generative distributions either collapse to a small set of outputs or become uniform over a large set, depending on the temperature setting. This finding is analogous to our observations with modulating CFG scales. Connecting our empirical analysis on finetuning Stable Diffusion with their theoretical framework could be an interesting future direction.
@@
Regarding mitigation strategies, existing works echo the importance of incorporating a large proportion of real data throughout the training loop [5, 12, 14] or accumulating more data [15]. The only mitigation strategy beyond altering the training data composition is proposed by [6], who suggests a self-correcting self-consuming loop using an expert model to correct synthetic outputs. While this approach is demonstrated in human motion generation with a physics simulator, having an expert model may not always be feasible for finetuning customized image generative models. In our work, we explore an alternative solution through reusable image generation. By making small improvements in finetuning and sampling steps, we can substantially enhance the reusability of images, allowing them to be used for many more iterations before the model starts to deteriorate. In a tangentially related work [9], the authors discuss bias amplification in self-consuming chains, noting minimal amplification over iterations.
3 Model Collapse in Self-Consuming Chain of Diffusion Finetuning
Table 1: Description of $10$ potential factors (excluding CFG) that we examined as candidate sources for model collapse. Dash (--) in table entries indicates that the default value was used. All experiments were conducted on Pokemon dataset except for the dataset size. For dataset size, we use CelebA since its original dataset is bigger than Pokemon and we can subsample $500$, $1000$, $2000$ images. More details in Appendix A.
Candidates Default 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18
Learning Rate 1e-4 1e-5 1e-3 -- -- -- -- -- -- -- -- -- -- -- -- -- -- -- --
Finetune Both -- -- Unet Text -- -- -- -- -- -- -- -- -- -- -- -- -- --
Noise No -- -- -- -- Yes -- -- -- -- -- -- -- -- -- -- -- -- --
Prompt [^1] C, W, S, and L for Combine, Waifu, Short, and Long, respectively. We concatenate BLIP and Waifu captions as default setting, referred as Combine. Short and Long prompts are BLIP captions generated with limitations in the lengths of captions. More details can be found in Appendix C.4. C -- -- -- -- -- W S L -- -- -- -- -- -- -- -- -- --
CLIP skip $2$- -- -- -- -- -- -- -- $1$ $3$- -- -- -- -- -- -- --
Diffusion Steps $30$- -- -- -- -- -- -- -- -- -- $100$- -- -- -- -- -- --
Epochs $100$- -- -- -- -- -- -- -- -- -- -- $50$ $200$- -- -- -- --
Real mixing $0.0$- -- -- -- -- -- -- -- -- -- -- -- -- $0.5$ $0.9$- -- --
Img / Prompt $1$- -- -- -- -- -- -- -- -- -- -- -- -- -- -- $5$- --
Dataset Size $1000$- -- -- -- -- -- -- -- -- -- -- -- -- -- -- -- $500$ $2000$
@@ 3.1 Problem setting & Experimental Setup
Chain of Diffusion
We begin with formally defining the self-consuming chain of diffusion finetuning. Given a pretrained generative model $M_{0}$, an original training image set $D_{0}=\{x_{0,i}|i\in[0,N-1]\}$, and a prompt set $P=\{y_{i}|i\in[0,N-1]\}$, where $N$ is the number of total images in the dataset, each image $x_{0,i}$ is paired with a corresponding text prompt $y_{i}$. $M_{k+1}$ is a model finetuned from $M_{0}$ using the generated image set $D_{k}=\{x_{k,i}|i\in[0,N-1]\}$ and the prompt set $P$, which simulates a common adaptation case of a pretrained generative model with a personally collected dataset. Then, $M_{k+1}$ generates a set of images $D_{k+1}$ for the next iteration using the prompt set $P$:
@@
(a) Reusability-FID trade-off [^3] We do not display scenarios that change the training dataset size, such as Img/Prompt and Dataset Size, as varying sizes result in FID scores on different scales. The related results are presented in the Appendix C, where we observe similar image degradation..
(b) Image degradation across different CFGs.
Figure 2: (a) Reusability-FID trade-off plot (Pokemon) shows that all scenarios other than changing CFG scales show poor reusability, amplifying the FID score $\sim$3x in 6 iterations of the chain. The x-axis represents $\text{FID}_{1}$, quantifying the generation performance at the first iteration and the y-axis represents the reusability defined in equation (4). For both $\text{FID}_{1}$ and reusability, lower is better. reusability (500 $\sim$ 1100) is about two times larger than $\text{FID}_{1}$ (300 $\sim$ 600), indicating $\text{FID}_{6}$ three times larger than $\text{FID}_{1}$. (b) Low CFG leads to blurry images and high CFG leads to high-frequency degradation in the Chains of Diffusion. CFG $2.5$ for Pokemon and $1.5$ for CelebA exhibit an ideal middle ground where both types of degradation slow down. More images in Appendix B.
@@
We use CLIP score [27] to quantify text-to-image alignment and Frechet Inception Distance (FID) score [28] to measure image fidelity. Furthermore, we compute the average sample-wise feature distance (SFD) between a pair of images corresponding to the text prompt as the fidelity metric applicable for text-to-image generation,
${\rm SFD}_{k}={\frac{1}{N}}\sum d(f(x_{k,i}),f(x_{0,i}))$ (3)
to evaluate each iteration $k$. SFD overcomes the problem of FID being sensitive to the number of images to compare. Following [29], we use DiNOv2 [30] as a feature extractor for both FID and SFD since it is more consistent with our visual inspection than Inception-V3 network [31]. With a slight abuse of terms, we will still refer to the Frechet distance with DiNOv2 features as the FID score.
New evaluation metric: Reusability.
In addition to evaluation metrics for generative models presented above, we propose a new metric to quantify the reusability of images generated in the Chain of Diffusion. We define a reusability metric as the performance discrepancy between the first iteration and the $K$-th iteration in the Chain of Diffusion:
$K\text{-reusability}={\rm FID}_{K}-{\rm FID}_{1},$ (4)
where ${\rm FID}_{k}$ stands for the FID between $k$-th iteration set and the original training set. We use FID as a performance metric here, but this can be CLIP, SFD, or any other performance metric of interest. Note that a low $K$-reusability indicates more reusable images since it means that the model does not degrade much in $K$ iterations. In the rest of the paper, $K=6$, and we simply call it reusability.
(a) t-SNE of images.
(b) CLIP.
(c) FID.
(d) SFD.
Figure 3: (a) t-SNE plot visualizes how the distribution of generated images shifts from the original distribution (black) in the Chain of Diffusion and shows distinct paths for three different CFG scales for Pokemon dataset. Different CFG scales and iterations are differentiated with different colors and transparency. Arrows indicate how the distributions of generated images of chains move for different CFG scales. While CFG 2.5 (blue) stays near the original images (black), high and low CFG scales (1.0 and 7.5) deviate fast, indicating image degradation. (b-d) Qualitative comparison for different Chains using CLIP $\uparrow$, FID $\downarrow$, and SFD $\downarrow$ for Pokemon dataset. CFG $2.5$ achieves the most robust performance for all three scores. CFG $1.0$ begins to gradually degrade from the beginning of the chain while CFG $7.5$ begins to degrade in the third iteration, which aligns with the visual inspection in Figure 22(b). Additional results for different datasets can be found in Appendix D
@@ 3.2 Model Collapse in the Chain of Diffusion
We observe significant image degradation in all four datasets (see Figure 11(b)) in the Chain of Diffusion. The quality of generated images begins to clearly deteriorate in the third iteration and it drops even more rapidly once the visible degradation emerges, reaching an unrecognizable level in two or three additional iterations. We investigated a variety of different scenarios (summarized in Table 1) to see if this degradation is an anomaly of specific hyperparameter settings or if it is a ubiquitous phenomenon. We tested various dataset sizes for $D_{0}$ and $D_{k}$ (Appendix C.1), increasing the size of synthetic datasets $D_{k}$ by generating more than one image per prompt (Appendix C.2), mixing real images from $D_{0}$ to $D_{k}$ (Appendix C.3), changing the descriptiveness of prompts (Appendix C.4), freezing U-Net or text encoder during finetuning (Appendix C.5), and various other hyperparameters (# sampling steps, # epochs, learning rate, and CLIP skip) (Appendix C.6, C.7, C.8, and C.9). We also tested adding a small Gaussian noise in each image in the original set $D_{0}$ to see if having small random perturbations can improve reusability (Appendix C.10), and investigated if the degradation happens for larger Stable Diffusion (Appendix C.11). In all settings we tested, image degradation was universally present. We plot the results for Pokemon dataset on Figure 23 where y-axis is reusability and x-axis is the ${\rm FID}_{1}$. While adding noise and mixing $90\%$ real images show the best reusability (the lower, the better), they still exhibit significant degradation (FID score has been doubled in $6$ iterations). More visual inspections and metric plots can be found in Appendix C.
Observation 2: CFG is the only factor that impacts the model collapse.
@@
High CFG means that we emphasize the conditional score for the given prompt more, which pushes the generation to align better with the prompt and often leads to higher-fidelity images. On the other hand, lower CFG places less weight on the conditional score and provides more diversity in generated images. For those familiar with temperature sampling [33], CFG plays a similar role as temperature, which adjusts the trade-off between fidelity and diversity.
In Figure 23, we observe that as we increase the CFG scale, the image quality in the first iteration improves (smaller ${\rm FID}_{1}$), which is expected from our understanding of CFG. A more surprising part is that this comes at the cost of a worse reusability metric (increase in ${\rm FID}_{6}-{\rm FID}_{1}$). Also, when the CFG scale is as high as $7.5$ or $10.0$, the improvement in ${\rm FID}_{1}$ plateaus, and increased CFG worsens both ${\rm FID}_{1}$ and reusability. Similarly, when the CFG scale is too low—below $2.0$—the improvement in reusability plateaus and both ${\rm FID}_{1}$ and reusability begin to decline. Interestingly, there is an optimal region of CFG values (near $2.5$, specific to Pokemon dataset), where we achieve good reusability while maintaining a good quality in the first iteration as well. Figure 3 presents quantitative results for Pokemon dataset to demonstrate how different CFG scales affect the performance in CLIP, FID, and SFD. CFG scale $2.5$ achieves the most robust performance for all metrics.
Figure 4: Latent analysis to investigate how different CFG scales affect the latent of Stable Diffusion (Pokemon dataset). (a) Latent distribution shrinks over iteration for low CFG and expands with high CFG. Larger values in latent vectors are more likely to occur with high CFG, gradually increasing the tail of the distribution. (b) Differences between conditional and unconditional scores increase as the training set is more degraded. Especially, high differences in the later diffusion steps can be a cause of high-frequency degradation.

### v2

# Model Collapse in the Self-Consuming Chain of Diffusion Finetuning: A Novel Perspective from Quantitative Trait Modeling

@@ Abstract
The success of generative models has reached a unique threshold where their outputs are indistinguishable from real data, leading to the inevitable contamination of future data collection pipelines with synthetic data. While their potential to generate infinite samples initially offers promise for reducing data collection costs and addressing challenges in data-scarce fields, the severe degradation in performance has been observed when iterative loops of training and generation occur—known as “model collapse.” This paper explores a practical scenario in which a pretrained text-to-image diffusion model is finetuned using synthetic images generated from a previous iteration, a process we refer to as the “Chain of Diffusion.” We first demonstrate the significant degradation in image quality caused by this iterative process and identify the key factor driving this decline through rigorous empirical investigations. Drawing an analogy between the Chain of Diffusion and biological evolution, we then introduce a novel theoretical analysis based on quantitative trait modeling. Our theoretical analysis aligns with empirical observations of the generated images in the Chain of Diffusion. Finally, we propose Reusable Diffusion Finetuning (ReDiFine), a simple yet effective strategy inspired by genetic mutations. ReDiFine mitigates model collapse without requiring any hyperparameter tuning, making it a plug-and-play solution for reusable image generation.
@@ 1 Introduction
Can state-of-the-art AI models learn from their own outputs and improve themselves? As generative AI models (e.g., GPT, Diffusion) now churn out uncountable synthetic texts and images, this question piqued the curiosity of many researchers in the past couple of years. While some show positive results of self-improving [1, 2], most report an undesirable “model collapse”—a phenomenon where a model’s performance degrades when it undergoes multiple cycles of training with the self-generated data [3, 4, 5, 6, 7, 8, 9, 10]. When large language models (LLMs) are trained with their own outputs, it begins to produce low-quality text with many repetitions [11], and its linguistic diversity declines rapidly [12, 13]; image models also show quality degradation [14, 15] and loss of diversity [16, 17].
The goal of this paper is to investigate model collapse in the practical scenario of finetuning pretrained text-to-image diffusion models. An end user of text-to-image diffusion models often wants to finetune the latest model to generate images with a very specific style (e.g., creating new characters in the style of Pokémon). In fact, hundreds of new finetuned diffusion models are regularly uploaded on platforms like CivitAI [^1] https://civitai.com/, each designed to produce different styles of images. When users scrape the internet to collect images in the style they want, it becomes almost inevitable that synthetic images will be included in their datasets. This is because the number of real images is limited, while synthetic images can be generated in massive quantities and dominate online sources. As a result, users will feel compelled to include more synthetic images in their datasets to meet the data demands of these ever-growing, data-hungry models.
To this end, we conduct a thorough investigation into how various hyperparameters commonly used during diffusion finetuning (e.g., learning rate, diffusion steps, prompts) impact model collapse. From our extensive empirical analysis, we make a crucial observation: the classifier-free guidance (CFG) scale is the most significant factor influencing the rate of model collapse. Moreover, we observe a fascinating phenomenon that the direction of degradation varies with different CFG values—low CFG results in low-frequency degradation in images, while high CFG leads to high-frequency degradation. The critical role of CFG in determining both the rate and direction of model collapse is a novel insight previously unrecognized in the literature.
To gain a deeper theoretical understanding of this phenomenon, we draw upon the concept of quantitative trait modeling from statistical genetics. Unlike existing studies that attribute model collapse to limited sample sizes and conclude that it leads to zero variance [16, 3, 6], we propose that the underlying cause is a truncation-based selection process modulated by CFG. This theoretical model accurately describes our experimental observations. Moreover, we show that other theoretical work on model collapse can also be connected to statistical genetics models, such as random genetic drift and its variants [18, 19, 20, 21]. This fresh angle of drawing parallels with statistical genetics offers a new framework to understand the mechanism of model collapse.
Finally, inspired by our empirical results and theoretical insights, we introduce the Reusable Diffusion Finetuning (ReDiFine) method. Our experiments show that selecting an appropriate CFG can significantly slow down model collapse. However, finding the optimal CFG is computationally expensive, requiring numerous iterations of finetuning. To address this, we propose ReDiFine as a plug-and-play solution that achieves a comparable reduction in collapse rate without the need for any hyperparameter tuning. The default ReDiFine setting effectively mitigates model collapse across all four datasets we tested. Moreover, ReDiFine is effective at mitigating model collapse not only in a worst-case scenario of having a fully synthetic dataset, but also in a more practical case, where we have a mix of real and synthetic images in the training set. Our contributions can be summarized as follows:
  • We performed thorough and extensive empirical investigations into model collapse when we finetune a diffusion model with diffusion-generated images. We tested an exhaustive list of experimental parameters on four datasets (two digital art & two natural images) to identify what affects model collapse. Our analysis reveals that CFG scale is the most influential factor that not only controls the rate of model collapse but also dictates the type of image degradation (Section 3).
  • We provide a novel theoretical analysis of model collapse based on quantitative trait modeling that can accurately predict how power spectra of generated images evolve over iterations (Section 4).
  • We propose a simple yet effective strategy to mitigate model collapse, named ReDiFine, which combines condition drop finetuning and CFG scheduling. Across all four datasets, ReDiFine successfully generates more “reusable” images, which can be reused to finetune a diffusion model without causing a severe model collapse (Section 5).
@@ 2 Related Work
The self-consuming training loop and the associated phenomenon known as “model collapse” have become significant areas of study in the past two years [15, 10, 5, 16, 14, 12, 3, 7, 13, 4, 8, 9]. Model collapse, defined as “a degenerative process affecting generations of learned generative models, where generated data end up polluting the training set of the next generation of models” in [6], has been observed in both language and image generative models.
Empirical studies on LLMs [13] reveal that linguistic diversity collapses, especially in high-entropy tasks [12], although this can be mitigated with data accumulation [2]. In image generation, several works [15, 10, 16, 17, 14, 3] note image degradation when diffusion models are recursively trained with self-generated data. We conduct extensive empirical experiments to identify the most significant factor causing model collapse in text-to-image diffusion models. Our findings reveal that while model collapse is universally observed across various datasets and scenarios, it manifests in three distinct types of image degradation.
(a) Self-Consuming Chain of Diffusion Finetuning.
(b) Image degradation in the Chain of Diffusion.
Figure 1: (a) Overall pipeline of the Chain of Diffusion. Given a pretrained text-to-image diffusion model $M_{0}$ and a prompt set $P$, a finetuned model $M_{k}$ is trained using $D_{k-1}$ generated at the previous iteration $k-1$. Then, $M_{k}$ generates $D_{k}$ using the same prompt set $P$, building a fully synthetic loop. Chain of diffusion begins with the original real dataset $D_{0}$. (b) Image degradation universally occurs during Chain of Diffusion across four datasets. As the Chain of Diffusion progresses, the severity of image degradation intensifies, exhibiting universal patterns of highly saturated images containing repetitive high-frequency patterns. This consistently holds across four datasets and $10$ scenarios that we comprehensively investigate in Section 3.2.
Theoretical studies on model collapse have largely focused on the reduction of diversity, typically framed as either decreasing covariance in continuous domains [16, 6, 3] or shrinking support in discrete domains [11, 9]. The finite number of generated samples per iteration has been identified as a key cause of model collapse by several authors [3, 6, 22, 11, 8], while others [16, 23, 9] emphasize that sampling bias reduces the generative model’s effective distribution. In contrast, we present a novel theoretical framework based on quantitative trait modeling, shifting the focus from variance reduction to mean drift induced by the selection process. We further demonstrate that many existing theories of model collapse can be understood as extensions of classical results from statistical genetics.
@@
Regarding mitigation strategies for model collapse, existing works echo the importance of incorporating a large proportion of real data throughout the training loop [16, 3, 8, 23] or accumulating additional data [2]. The only mitigation strategy beyond altering the training data composition is proposed by [4], who suggests a self-correcting self-consuming loop using an expert model to correct synthetic outputs. While this approach is demonstrated in human motion generation with a physics simulator, having an expert model may not be feasible and could be too costly for many real-world applications. In our work, we propose an alternative solution through reusable image generation, and show that it can mitigate model collapse more effectively than mixing more real data.
3 Model Collapse in Self-Consuming Chain of Diffusion
@@ 3.1 Problem Setting & Experimental Setup
Chain of Diffusion.
We begin with formally defining the self-consuming Chain of Diffusion finetuning. Given a pretrained generative model $M_{0}$, an original training image set $D_{0}=\{x_{0,i}|i\in[0,N-1]\}$, and a prompt set $P=\{y_{i}|i\in[0,N-1]\}$, where $N$ is the number of total images in the dataset, each image $x_{0,i}$ is paired with a corresponding text prompt $y_{i}$. $M_{k+1}$ is a model finetuned from $M_{0}$ using the generated image set $D_{k}=\{x_{k,i}|i\in[0,N-1]\}$ and the prompt set $P$, which simulates a fully synthetic loop [16] for finetuning a pretrained generative model. Then, $M_{k+1}$ generates a set of images $D_{k+1}$ for the next iteration using the prompt set $P$:
@@
Figure 2: Left: Description of $10$ potential factors (excluding CFG) that we examined as candidate sources for model collapse. All experiments were conducted on Pokemon except for the dataset size. For dataset size, we use CelebA since its original dataset is bigger, and we can subsample $500$, $1000$, and $2000$ images. For prompt, we concatenate prompts with different lengths.44footnotemark: 4 Right: All hyperparameter settings other than changing CFG show a high collapse rate greater than $100$, i.e., FID score increases by $\sim$2x in 6 iterations, indicating severe image degradation. The x-axis represents $\text{FID}_{1}$, quantifying the generation performance at the first iteration and the y-axis represents the collapse rate defined in Equation (3). For both $\text{FID}_{1}$ and collapse rate, lower is better.55footnotemark: 5
@@
We use Frechet Inception Distance (FID) [31] to measure image fidelity. Following [32], we use DiNOv2 [33] as a feature extractor for FID since it is more consistent with our visual inspection than Inception-V3 network [34]. With a slight abuse of terms, we will still refer to the Frechet distance with DiNOv2 features as the FID score. Quantitative results for additional metrics including CLIP score [35] can be found in Appendix E.
In addition to evaluation metrics for generative models presented above, we propose a new metric to quantify the reusability of images generated in the Chain of Diffusion. We define collapse rate as the performance degeneration per iteration in the Chain of Diffusion:
collapse $\text{collapse rate}=\Delta{\rm FID}=\frac{{\rm FID}_{K}-{\rm FID}_{1}}{K-1},$ (3)
where ${\rm FID}_{k}$ stands for the FID between $k$-th iteration set and the original training set. We use FID as a performance metric here, but this can be CLIP, or any other performance metric of interest. Note that a low collapse rate indicates more reusable images since it means that the model does not degrade much during the Chain of Diffusion and we have $K=6$ in the rest of the paper.
@@ 3.2 Model Collapse in the Chain of Diffusion
We observe significant image degradation in all four datasets (see Figure 11(b)) in the Chain of Diffusion. The quality of generated images begins to clearly deteriorate in the third iteration and drops even more rapidly once the visible degradation emerges, reaching an unrecognizable level in two or three additional iterations. We investigated a variety of different scenarios (summarized in Figure 2) to see if this degradation is an anomaly of specific hyperparameter settings or if it is a ubiquitous phenomenon. We tested various dataset sizes for $D_{0}$ and $D_{k}$, increasing the size of synthetic datasets $D_{k}$ by generating more than one image per prompt, mixing real images from $D_{0}$ to $D_{k}$, changing the descriptiveness of prompts, freezing U-Net or text encoder during finetuning, and various other hyperparameters (# sampling steps, # epochs, learning rate, and CLIP skip). We also tested adding a small Gaussian noise in each image in the original set $D_{0}$ to see if having small random perturbations can improve reusability, and investigated if the degradation happens for larger Stable Diffusion. In all settings we tested, image degradation was universally present and very fast. We plot the results for Pokemon on Figure 2 where y-axis is collapse rate and x-axis is the ${\rm FID}_{1}$ (better when closer to the origin). While adding noise to images and mixing $90\%$ real images to every iteration as proposed by [16, 3, 8, 23, 2] show the lowest collapse rate, they still exhibit significant degradation as shown in quantitative values (FID score has been doubled in $6$ iterations). Visual inspection for images is provided in Appendix C.
Observation 2: CFG is the most significant factor that impacts the model collapse.
@@
A higher CFG means that we emphasize the conditional score for the given prompt more, which pushes the generation to align better with the prompt and often leads to higher-fidelity images. On the other hand, lower CFG places less weight on the conditional score and provides more diversity in generated images. For those familiar with temperature sampling [37], CFG plays a similar role as temperature, which adjusts the trade-off between fidelity and diversity.
In Figure 2, we observe that as we increase the CFG scale, the image quality in the first iteration improves (smaller ${\rm FID}_{1}$ on x-axis), which is expected from our understanding of CFG. A more surprising part is that this comes at the cost of a worse collapse rate (an increase on the y-axis). Additionally, when the CFG scale is as high as $7.5$ or $10.0$, the improvement in ${\rm FID}_{1}$ plateaus, and increased CFG worsens both ${\rm FID}_{1}$ and collapse rate. Similarly, when the CFG scale is too low—below $2.0$—the improvement in the collapse rate plateaus and both ${\rm FID}_{1}$ and collapse rate begin to increase. There is an optimal region of CFG values (near $2.5$, specific to Pokemon), where we achieve low collapse rate while maintaining a good quality in the first iteration as well. Moreover, Figure 33(c) presents FID for Pokemon to demonstrate how different CFG scales affect the performance during iterations. CFG scale $2.5$ achieves the most robust performance for Pokemon across iterations. Interestingly, optimal CFG scales differ for different styles: $2.5$ for animated or hand-writing datasets (Pokemon and Kumapi) and $1.5$ for photo datasets (CelebA and Butterfly) as shown in Appendix B.
(a) Degradations across different CFGs.
(b) t-SNE of images.
(c) FID $\downarrow$.
Figure 3: (a) Low CFG leads to blurry images and high CFG leads to high-frequency degradation in the Chain of Diffusion. CFG $2.5$ for Pokemon and $1.5$ for CelebA exhibit an ideal middle ground where both types of degradation slow down. More images in Appendix B. (b) t-SNE plot visualizes how generated images evolve from the original distribution (black) and shows distinct paths for three CFG scales (Pokemon). Different CFG scales and iterations are differentiated with different colors and transparency. Arrows indicate how the distributions of generated images move for different CFG scales. While CFG 2.5 (blue) stays near the original images (black), high and low CFG scales (7.5 and 1.0) deviate fast, indicating image degradation. (c) Quantitative comparison for FID $\downarrow$ (Pokemon). CFG $2.5$ achieves the most robust performance. CFG $1.0$ degrades from the beginning of the chain while CFG $7.5$ begins to degrade in the third iteration, which aligns with the visual inspection in (a).